Find the distance between points P1=(3,-1,5)P1=(3,-1,5) and P2=(2.1.-1). Find the distance between points P1=(1.-5,4)P1=(1.-5Please answer all using MATLAB
Answer #1

a)

>> P1=[3;-1;5]

P1 =

     3
    -1
     5

>> P2=[2;1;-1]

P2 =

     2
     1
    -1

>> norm(P1-P2,2)

ans =

    6.4031

b)

>> P1=[1;-5;4]

P1 =

     1
    -5
     4

>> P2=[4;-1;-1]

P2 =

     4
    -1
    -1

>> norm(P1-P2,2)

ans =

    7.0711

c)

c=[-4,-9]

c =

    -4    -9

>> d=[-1,2]

d =

    -1     2

>> d=dot(c,d)


d =

   -14

d)

>> c=[2 1 -1]

c =

     2     1    -1

>> d=[1 2 0]

d =

     1     2     0

>> d=dot(c,d)

d =

     4

e)

>> mu=5

mu =

     5

>> syms x
>> T = taylor((1+x)^mu)
 
T =
 
x^5 + 5*x^4 + 10*x^3 + 10*x^2 + 5*x + 1
 

I have shown matlab command for \mu =5 .

f)

>> syms x
>> f = x^3;
>> T=taylor(f, x,'ExpansionPoint',3, 'Order', 8)
 
T =
 
27*x + 9*(x - 3)^2 + (x - 3)^3 - 54